List all real values of x such that f(x) = 0.
f(x) = {-6 x + 11}/{-18}

Answer by stanbon(75887)

List all real values of x such that f(x) = 0.
f(x) = {-6 x + 11}/{-18}
---------
(-6x + 11)/(-18) = 0
-6x + 11 = 0
6x = 11
x = 11/6
===================
Cheers,
Stan H.